1. Suppose that ABC Corp. has an equity beta of 1.5. Current risk-free rates are 5% and the expected return on the market portfolio is 15%. ABC Corp has a market value debt-to-equity ratio of 0.5, debt that is rated AA, and faces a 21% tax rate. Assuming that the credit spread on AA debt is 2%, compute ABC Corps WACC? Enter your answer as a percent without the % sign. Round to two decimals.

2. Suppose that QRY Corp. has an equity beta of 1.5. Current risk-free rates are 5% and the expected return on the market portfolio is 15%. QRY Corp has a market value debt-to-equity ratio of 0.5, debt that is rated AAA, and faces a 21% tax rate. QRY Corps debt is rate AAA and has an average maturity of 10 years. Assuming that the current market price of AAA bonds paying 5% semi-annual coupon, with 10 years maturity, and par value of $1000 is $850, compute QRY Corps WACC? Enter your answer as a percent without the % sign. Round to two decimals.
Answer #1

As per HOMEWORKLIB RULES, only one question is answered.

Question 1:

Given information:

Equity beta = 1.5%

risk free rate = Rf = 5%

Market return = Rm = 15%

Debt to Equity ration = D/E = 0.5 hence D = 0.5 * E

Tax rate = 21%

Since the credit spread is 2%, cost of debt = Rd = risk free rate + credit spread = 5 + 2 = 7%

WACC is given by the formula:

WACC = E/(D+E) * cost of equity + D/(D+E) * cost of debt

we will calculate cost of equity (Re) using CAPM model:

cost of equity = Re = beta * (Market return - risk free rate) + risk free rate

= 1.5 *(15 - 5) + 5 = 20 %

cost of debt = interest rate * (1- tax rate) = 7% * (1-21%) = 5.53%

substitute cost of equity and cost of debt in the WACC formula:

WACC = E/(D+E) * cost of equity + D/(D+E) * cost of debt

= E/(0.5E +E) * 20% + 0.5E/(0.5E+E) * 5.53%

= E/1.5E *20% +0.5E/1.5E *5.53%

= 1/1.5 *20% + 0.5/1.5 * 5.53% = 0.1518 = 15.18%

Answer: WACC = 15.18%